OpenStudy (anonymous):

please help @Smileyside and @cherry18

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

k

OpenStudy (anonymous):

what is \[\sqrt[3]{-216}\]

OpenStudy (anonymous):

the decimal form is 5.9999999, round it and you will get your answer

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what is \[\sqrt[3]{72}\]

OpenStudy (anonymous):

i got 4.1601676461

OpenStudy (anonymous):

well here are the choices A. \[9\sqrt[3]{2}\]

OpenStudy (anonymous):

B. \[3\sqrt[3]{9}\]

OpenStudy (anonymous):

C. \[2\sqrt[3]{9}\]

OpenStudy (anonymous):

and D. \[9\sqrt[3]{3}\]

OpenStudy (anonymous):

i think c because it had the same decimal point as the question

OpenStudy (anonymous):

i have to go now

OpenStudy (anonymous):

@welshfella

OpenStudy (ahvei):

its C

OpenStudy (anonymous):

are you sure?

OpenStudy (ahvei):

yes

OpenStudy (anonymous):

C. and sorry I was late had softball practice.